EE210 HW i] write the Mesh 42 2] write the Mesh Egs 3] Write the Mesh Egs CA © ¢6 Is <8
Answer #1

at loop o, 1 -10+29, +6 (1,-1) = 0 at Coop @ , 1 -6(1-1) + 3 I₂ +12=0 at loop ③ , 1.-12 +423 +5 Iz=0 SVODE dat loop. a. -6+6(Y at loop o -6( Ly - 60) + 4 ( [x - 5y) +81470- I at Coop. - Fly 4 (Ix-Iy) = 0.